# Multispectral State-Space Feature Fusion: Bridging Shared and Cross-Parametric Interactions for Object Detection

### v1

@@ 4.2 Experimental Setup
4.3 Ablation Studies
4.3.1 Different Detection Frameworks
To validate the generalizability of the proposed MS2Fusion module, we conducted experiments using two mainstream detection frameworks, YOLOv5 and CoDetr [38]. In both frameworks, add fusion are employed as the baseline method for comparison. The results in Table 2 demonstrate that our fusion method achieves significant performance improvement over the baseline model in both the YOLOv5 and CoDetr frameworks. These experimental results fully verify the effectiveness and broad applicability of MS2Fusion, confirming that it serves as a plug-and-play universal module that can be flexibly integrated into different types of detection frameworks.
Table 2: Effect of different backbones and detection frameworks.
Num Framework Fusion Model Person Car Bicycle mAP@0.5 FPS
1 YOLOv5 Baseline 83.8 88.9 67.3 80.0 43.2
2 MS2Fusion 85.1 89.8 74.9 83.3 (+3.3) 24.4
3 CoDetr Baseline 88.0 91.7 73.8 84.5 5.9
4 MS2Fusion 90.2 93.4 79.6 87.8 (+3.3) 4.8
4.3.2 Different Backbones
We have conducted experiments using three different backbones: VGG16 [44], ResNet50 [45], and CSPDarkNet53 [16]. As shown in Table 3, the performance of MS2Fusion has achieved a consistent improvement in all settings. Compared to the baseline, our method improves 1.0%, 4.8%, and 3.3% with the backbone of VGG16, ResNet50 and CSPDarkNet53, respectively. These results also show that the MS2Fusion module is not only effective but also has good generalization to different backbone networks.
Table 3: Effect of different backbones.
Num Backbone Fusion Model Person Car Bicycle mAP@0.5
1 VGG16 Baseline 78.9 87.7 51.2 72.6
2 MS2Fusion 79.0 87.8 53.8 73.6 (+1.0)
3 ResNet50 Baseline 76.8 85.7 44.6 69.0
4 MS2Fusion 80.2 88.4 52.8 73.8 (+4.8)
5 CSPDarkNet53 Baseline 83.8 88.9 67.3 80.0
6 MS2Fusion 85.1 89.8 74.9 83.3 (+3.3)
Table 4: Effect of different fusion modules.
Num CP-SSM SP-SSM FF-SSM Person Car Bicycle mAP@0.5 Params(M)
1 83.8 88.9 67.3 80.0 72.7
2 83.8 89.8 72.7 82.1 (+2.1) 84.5
3 85.6 90.4 70.6 82.2 (+2.2) 90.4
4 84.9 90.1 72.4 82.5 (+2.5) 86.0
5 84.9 90.3 72.9 82.7 (+2.7) 117.0
6 85.5 90.0 72.5 82.7 (+2.7) 97.8
7 85.5 90.3 72.9 82.9 (+2.9) 103.7
8 85.1 89.8 74.9 83.3 (+3.3) 130.2
4.3.3 Different Modules
Our comparative analysis of different feature fusion modules (CP-SSM, SP-SSM, and FF-SSM) in Table 4 reveals their distinct strengths in object detection tasks. Furthermore, the results also highlight the synergistic benefits achieved through their integration. The experimental results demonstrate that CP-SSM, as a modality-specific feature enhancement module, brings significant performance gains for challenging scenarios. Specifically, it achieves a 2.1% mAP improvement when deployed independently, with particularly notable gains of 5.4% on the ’bicycle’ category, effectively addressing the small object detection challenge. In contrast, SP-SSM specializes in cross-modal shared feature extraction, exhibiting superior performance on structurally well-defined categories such as ’cars’ while maintaining parameter efficiency. The shared representation learned by SP-SSM demonstrates strong generalization across modalities. Besides, the FF-SSM module establishes an efficient framework for heterogeneous feature fusion, achieving a 2.5% mAP boost with fewer computational overhead. More importantly, FF-SSM maintains robust performance across all object categories, demonstrating its effectiveness as a unified fusion solution.
Further analysis reveals significant synergistic effects when combining these modules. The joint use of SP-SSM and FF-SSM is especially outstanding (mAP 82.9%), demonstrating that shared features optimized by the fusion module exhibited stronger performance. The combination of all three modules achieves the best performance (mAP 83.3%), with bicycle detection accuracy improving by 7.6% compared to the baseline. This is attributed to the fine-grained features provided by CP-SSM, the common patterns extracted by SP-SSM, and the dynamic feature fusion enabled by FF-SSM. Notably, the increase in parameter number for the three-module combination is proportionally reasonable relative to the performance gains, proving the efficiency of this design.
In summary, the proposed CP-SSM, SP-SSM, and FF-SSM modules collectively enhance model performance through three synergistic mechanisms: feature complementarity, feature sharing and feature fusion. The hierarchical architecture effectively preserves category-specific distinctive features while discovering cross-category common patterns, with adaptive fusion eliminating information redundancy. This fusion approach provides an effective solution for object detection in complex-scenario that achieves an optimal accuracy-efficiency balance, offering three key insights for multi-modal feature fusion network design: the importance of modality-specific feature enhancement for challenging categories, the benefits of cross-modal shared representations for improved generalization and the effectiveness of adaptive fusion mechanisms in maximizing complementary benefits while minimizing redundancy.
4.3.4 Different Fusion Layers
As shown in Table 5, the incremental addition of MS2Fusion modules across different feature stages yields distinct performance characteristics. Initial test at only the P3 stage (Row 2) demonstrates a trade-off effect, where ’bicycle’ AP experiences a substantial 5.7% improvement to 71.4% at the cost of marginal decreases in ’person’ (83.4%) and ’car’ (88.5%) detection, ultimately elevating the mAP to 81.1%. Expanding the fusion to both P3 and P4 stages (Row 3) reverses this pattern, boosting ’person’ and ’car’ AP to 85.8% and 90.1%. When employing fusion at all three stages (Row 4), it achieves optimal balance. While maintaining strong performance on ’person’ (85.1%) and ’car’ (89.8%), it delivers a remarkable 74.9% AP for ’bicycles’, the highest among all configurations and pushes the overall mAP to 83.3%. This systematic evaluation clearly demonstrates that multi-level feature fusion enables more effective cross-modal feature integration, with full-stage implementation proving particularly advantageous for challenging categories like ’bicycles’ while preserving performance on other objects.
Table 5: Effect of fusion at different layers.
Number P3 P4 P5 Person Car Bicycle mAP@0.5
1 83.8 88.9 67.3 80.0
2 83.4 88.5 71.4 81.1 (+1.1)
3 85.8 90.1 70.0 82.0 (+2.0)
4 85.1 89.8 74.9 83.3 (+3.3)
Table 6: Finetunes in CP-SSM.
Num Finetune Person Car Bicycle mAP@0.5
1 rows 85.1 89.8 74.9 83.3
2 columns 64.1 90.3 72.8 82.4
3 rows and columns 85.4 90.4 70.2 82.0
4 w/o exchange $C$ 85.7 90.6 71.7 82.7
5 exchange $C$ 85.1 89.8 74.9 83.3
4.3.5 Discussion on CP-SSM
Exchange $C$-parameters. As shown in Table 6 (rows 4-5), our investigation of cross-branch C-parameter exchange within the CP-SSM module reveals significant performance benefits. The proposed parameter sharing mechanism yields a consistent 0.6% performance gain compared to non-exchanging configurations, establishing an effective implicit enhancement through shared hidden state projection matrices. This innovative design achieves three key advantages: (1) enhanced joint feature capture capability, particularly for complex data relationships; (2) improved cross-feature interaction through optimized projection matrix sharing; and (3) strengthened representation learning via complementary information flow. Experimental validation confirms that this parameter exchange strategy is particularly effective in scenarios requiring sophisticated data representation, with the CP-SSM module demonstrating superior performance in cross-modal feature extraction tasks. The success of this approach reveals the importance of carefully designed parameter interaction mechanisms in modern feature learning architectures.
4.3.6 Comparison with Different Input Modalities
Num Model Input mAP@0.5
1 YOLOv5 V 67.8
2 T 73.9
3 Baseline V+V 61.2
4 T+T 77.8
5 V+T 80.0
6 Ours V+V 68.4 (+7.2)
7 T+T 82.1 (+4.3)
8 V+T 83.3 (+3.3)
Table 8: Input Configuration for FF-SSM Module
FF-SSM input mAP@0.5 person car bicycle
(V, T) 83.1 85.4 90.2 73.9
(T, V) 82.5 85.1 90.1 72.4
((V, T), (T, V)) 83.3 85.1 89.8 74.9
(a) (V, T))
(b) (T, V)
(c) ((V, T), (T, V))
4.3.7 Analysis of FF-SSM Input Configurations
In this section, we have investigated the effect of the FF-SSM module with three different input order in SSM: (V, T), (T, V), and bidirectional input ((V, T), (T, V)). LABEL:FF-SSM_input demonstrates significant performance variations with single-order inputs: the (V, T) configuration achieved 83.1% (0.2% drop) on mAP@0.5, with particularly notable degradation in ’bicycle’ detection accuracy (1.0% drop), while (T, V) decreased to 82.5% (0.8% drop), with particularly notable degradation in ’bicycle’ detection accuracy (2.5% drop). This reveals a feature attenuation phenomenon where the model gradually forgets early input features during state propagation. To address this, our proposed bidirectional architecture establishes cross-modal feature memory pathways. This solution not only improves overall performance to 83.3% but also significantly enhances ’bicycle’ detection accuracy (74.9%), while maintaining detection precision for ’person’ and ’car’.
As shown in Figure 10, we compared three different input configurations of ERF maps for the FF-SSM module: (V, T), (T, V) and ((V, T), (T, V)). It clearly demonstrates that our method achieves a significantly larger receptive field when using the bidirectional input strategy compared to the unidirectional input modes. This experimental result strongly validates the effectiveness of our proposed bidirectional input architecture in expanding the model’s receptive field. The design effectively resolves feature attenuation in state-space models at the cost of linear computational complexity, offering a novel solution for cross-modal fusion.
@@ 4.4.1 Comparison on the FLIR Dataset
[…]
@@ 4.5.1 Experiments on RGB-T Semantic Segmentation
MFNet dataset: It is the first publicly available RGB-T dataset featuring pixel-level annotations. It consists of 1,569 aligned RGB-T image pairs captured in urban environments, with semantic labels for eight common driving-scene obstacles: cars, pedestrians, bicycles, curves, bus stops, guardrails, traffic cones, and speed bumps.
@@ 4.5.2 Experiments on RGB-T Salient Object Detection (RGB-T SOD)
Evaluation metrics: Mean Absolute Error (MAE) measures the average magnitude of errors between prediction and ground truth, without considering their direction.
The F-Measure (adpF) is designed to assess classification performance, especially in situations with imbalanced class distributions. It extends the traditional F-measure (F1-score) by allowing adjustments to better handle varying levels of class imbalance or different importance of precision and recall. S-Measure (S) is a metric used to evaluate the performance of image segmentation models by assessing both boundary accuracy and region consistency. It provides a comprehensive measure of segmentation quality by incorporating the structure of the segmentation and its alignment with the ground truth. E-Measure (adpE) adapts to different conditions by an adaptive threshold that is set to twice the mean values of the salient maps.
@@ 4.6.1 Heatmap Visualization
Figure 11 presents the heatmap comparisons of three competing methods (baseline, ICAFusion and MS2Fusion) on RGB-T images. Through comprehensive analysis across multiple scenarios, the MS2Fusion method demonstrates consistent superiority. In the parking lot scenario (first row), the baseline method manages to detect partial instances of cars and pedestrians, yet produces significantly incomplete bounding boxes. Although ICAFusion shows noticeable improvement by capturing more objects, it still suffers from occasional missed detections. By contrast, MS2Fusion achieves near-perfect performance, precisely localizing all cars and pedestrians with highly accurate bounding boxes. The performance gap becomes even more pronounced in the challenging nighttime street scene (second row). While the baseline method can identify some pedestrians and vehicles, its detection accuracy proves inadequate. ICAFusion offers moderate improvement over the baseline, yet still struggles with low detection accuracy in these low-light conditions. Remarkably, MS2Fusion maintains excellent performance, reliably identifying all targets with exceptional precision even in this demanding scenario. Similarly, in the road scene (third row), the baseline method exhibits several limitations, including missed detections of vehicles and cyclists. ICAFusion partially addresses these issues by detecting more instances, but its accuracy remains suboptimal. Impressively, MS2Fusion again outperforms both competitors, achieving better detection performance with precise bounding boxes and minimal false negatives.

### v2

@@ 4.3.1 Comparison on the FLIR Dataset
Table 2: Comparison on the FLIR-align dataset. (’-’ indicates missing values. $\ddagger$}’ symbol denotes experimental results obtained using the CoDetr framework, with input images resized to a fixed resolution of 640×640 pixels.)
@@
Fusion-Mamba(yolov5) [60] 84.3 44.4 - - -
Fusion-Mamba(yolov8) [60] 84.9 47.0 - - -
@@
CPCF [32] 96.4 65.0
Fusion-Mamba(yolov5) [60] 96.8 62.8
Fusion-Mamba(yolov8) [60] 97.0 64.3
@@ 4.3.2 Comparison on the LLVIP Dataset
Table 5: Comparison on the VEDAI dataset.
Methods mAP@0.5 Car Truck Pickup Tractor Camper Ship Van Plane
ICAFusion [8] 76.6 88.0 67.5 80.9 74.9 78.3 56.9 69.8 96.7
MidFusion1 [48] 77.4 88.2 84.9 81.7 68.4 72.2 71.0 53.6 99.5
MidFusion2 [48] 78.9 89.2 78.8 87.6 66.2 73.2 62.5 73.9 99.5
Input Fusion1 [48] 80.0 88.7 77.7 79.5 72.3 75.7 78.6 72.1 95.7
Input Fusion2 [48] 80.5 89.1 84.9 78.7 84.2 70.1 69.9 68.8 98.4
YOLOFusion [48] 81.6 91.7 78.1 85.9 71.9 78.9 71.7 75.2 99.5
Ours 80.2 89.5 75.6 84.2 81.6 75.9 61.8 76.4 97.0
Ours‡ 84.2 93.3 78.9 88.5 85.8 80.6 65.5 83.6 97.1
@@ 4.3.3 Comparison on the M3FD Dataset
4.3.4 Comparison on the VEDAI dataset.
Table 5 highlights that our method achieves state-of-the-art performance on the VEDAI dataset, even without specialized techniques for detecting small objects. This underscores our model’s strength in focusing on the multi-modal features of small objects, proving its effectiveness. These results indicate that our approach excels in UAV image applications, ensuring robust performance in detecting small objects.
@@ 4.4.1 Experiments on RGB-T Semantic Segmentation
MFNet dataset [72]: It is the first publicly available RGB-T dataset featuring pixel-level annotations. It consists of 1,569 aligned RGB-T image pairs captured in urban environments, with semantic labels for eight common driving-scene obstacles: cars, pedestrians, bicycles, curves, bus stops, guardrails, traffic cones, and speed bumps.
@@ 4.4.2 Experiments on RGB-T Salient Object Detection (RGB-T SOD)
Evaluation metrics: Following [97], Mean Absolute Error (MAE), F-Measure (adpF), S-Measure (S), and E-Measure (adpE) are commonly used as evaluation metrics for SOD.
@@
4.5 Ablation Studies
4.5.1 Different Detection Frameworks
To thoroughly evaluate the generalization of the MS2Fusion module, comparative experiments are conducted on two mainstream detection frameworks, YOLOv5 and CoDetr [41], with add fusion serving as the baseline method. Table 9 unequivocally demonstrate that our fusion approach achieves remarkable performance improvements in both frameworks, delivering a 3.3% boost in mAP@0.5 over the baseline. These findings strongly validate the efficacy and versatility of MS2Fusion. Notably, the module exhibits excellent plug-and-play characteristics, enabling seamless integration with various detection frameworks and providing a flexible yet powerful solution for multimodal object detection tasks.
Table 9: Effect of different backbones and detection frameworks.
Num Framework Fusion Model Person Car Bicycle mAP@0.5 FPS
1 YOLOv5 Baseline 83.8 88.9 67.3 80.0 43.2
2 MS2Fusion 85.1 89.8 74.9 83.3 (+3.3) 24.4
3 CoDetr Baseline 88.0 91.7 73.8 84.5 5.9
4 MS2Fusion 90.2 93.4 79.6 87.8 (+3.3) 4.8
4.5.2 Different Backbones
To comprehensively evaluate the architectural compatibility of the MS2Fusion module, we conducted systematic experiments with three distinct backbone networks: VGG16 [98], ResNet50 [99], and CSPDarkNet53 [18]. As demonstrated in Table 10, MS2Fusion achieves consistent performance gains across all architectural configurations. Specifically, our method yields improvements of 1.0%, 4.8%, and 3.3% over the baseline when employing VGG16, ResNet50, and CSPDarkNet53 backbones, respectively. These experimental results not only validate the effectiveness of the MS2Fusion module but also substantiate its remarkable generalization capability across diverse backbone architectures.
Table 10: Effect of different backbones.
Num Backbone Fusion Model Person Car Bicycle mAP@0.5
1 VGG16 Baseline 78.9 87.7 51.2 72.6
2 MS2Fusion 79.0 87.8 53.8 73.6 (+1.0)
3 ResNet50 Baseline 76.8 85.7 44.6 69.0
4 MS2Fusion 80.2 88.4 52.8 73.8 (+4.8)
5 CSPDarkNet53 Baseline 83.8 88.9 67.3 80.0
6 MS2Fusion 85.1 89.8 74.9 83.3 (+3.3)
Table 11: Effect of different fusion modules.
Num CP-SSM SP-SSM $F$-SSM Person Car Bicycle mAP@0.5 Params(M)
1 83.8 88.9 67.3 80.0 72.7
2 83.8 89.8 72.7 82.1 (+2.1) 84.5
3 85.6 90.4 70.6 82.2 (+2.2) 90.4
4 84.9 90.1 72.4 82.5 (+2.5) 86.0
5 84.9 90.3 72.9 82.7 (+2.7) 117.0
6 85.5 90.0 72.5 82.7 (+2.7) 97.8
7 85.5 90.3 72.9 82.9 (+2.9) 103.7
8 85.1 89.8 74.9 83.3 (+3.3) 130.2
4.5.3 Different Modules
Our comparative study of CP-SSM, SP-SSM, and FF-SSM modules reveals their complementary strengths in object detection. CP-SSM excels in enhancing modality-specific features, improving mAP by 2.1% and boosting challenging category detection (e.g., bicycles) by 5.4%. SP-SSM focuses on extracting cross-modal shared features with strong generalization, while FF-SSM delivers efficient heterogeneous feature fusion, yielding a 2.5% mAP gain with minimal overhead and consistent performance across categories.
Combining modules further amplifies performance. SP-SSM + FF-SSM achieves 82.9% mAP, while integrating all three modules pushes mAP to 83.3%, improving bicycle detection by 7.6%. This synergy balances feature complementarity, sharing, and fusion, enhancing accuracy while keeping parameter growth reasonable.
In summary, CP-SSM, SP-SSM, and FF-SSM work in concert to boost detection via modality-specific enhancement, cross-modal generalization, and adaptive fusion—preserving unique features, uncovering shared semantics, and reducing redundancy effectively.
4.5.4 Different Fusion Layers
As shown in Table 12, the incremental addition of MS2Fusion modules across different feature stages yields distinct performance characteristics. Initial test at only the P3 stage (Row 2) demonstrates a trade-off effect, where ’bicycle’ AP experiences a substantial 5.7% improvement to 71.4% at the cost of marginal decreases in ’person’ (83.4%) and ’car’ (88.5%) detection, ultimately elevating the mAP to 81.1%. Expanding the fusion to both P3 and P4 stages (Row 3) reverses this pattern, boosting ’person’ and ’car’ AP to 85.8% and 90.1% When employing fusion at all three stages (Row 4), it achieves optimal balance. While maintaining strong performance on ’person’ (85.1%) and ’car’ (89.8%), it delivers a remarkable 74.9% AP for ’bicycles’, the highest among all configurations and pushes the overall mAP to 83.3%. This systematic evaluation clearly demonstrates that multi-level feature fusion enables more effective cross-modal feature integration, with full-stage implementation proving particularly advantageous for challenging categories like ’bicycles’ while preserving performance on other objects.
Table 12: Effect of fusion at different layers.
Number P3 P4 P5 Person Car Bicycle mAP@0.5
1 83.8 88.9 67.3 80.0
2 83.4 88.5 71.4 81.1 (+1.1)
3 85.8 90.1 70.0 82.0 (+2.0)
4 85.1 89.8 74.9 83.3 (+3.3)
Table 13: Finetunes in CP-SSM.
Num Finetune Person Car Bicycle mAP@0.5
1 rows 85.1 89.8 74.9 83.3
2 columns 64.1 90.3 72.8 82.4
3 rows and columns 85.4 90.4 70.2 82.0
4 w/o exchange $C$ 85.7 90.6 71.7 82.7
5 exchange $C$ 85.1 89.8 74.9 83.3
4.5.5 Discussion on CP-SSM
Exchange $C$-parameters. As shown in Table 13, the experimental study of the cross-branch C-parameter exchange mechanism in the CP-SSM module demonstrates that exchanging hidden state projection matrices can significantly improve model performance. Compared to non-exchange configurations, this mechanism achieves a stable performance improvement of 0.6%, not only enhancing the feature capture capability for complex data relationships but also optimizing cross-feature interaction efficiency. Experimental validation shows that this parameter exchange strategy is particularly effective in scenarios requiring fine-grained data representation, enabling the CP-SSM module to demonstrate outstanding performance in cross-modal feature extraction tasks. This finding highlights the importance of parameter interaction mechanism design in modern feature learning architectures.
4.5.6 Comparison with Different Input Modalities
Num Model Input mAP@0.5
1 YOLOv5 V 67.8
2 T 73.9
3 Baseline V+V 61.2
4 T+T 77.8
5 V+T 80.0
6 Ours V+V 68.4 (+7.2)
7 T+T 82.1 (+4.3)
8 V+T 83.3 (+3.3)
Table 15: Input Configuration for FF-SSM Module
FF-SSM input mAP@0.5 person car bicycle
(V, T) 83.1 85.4 90.2 73.9
(T, V) 82.5 85.1 90.1 72.4
((V, T), (T, V)) 83.3 85.1 89.8 74.9
(a) (V, T))
(b) (T, V)
(c) ((V, T), (T, V))
4.5.7 Analysis of FF-SSM Input Configurations
In this section, we systematically evaluate the FF-SSM module with three distinct input configurations: (V, T), (T, V), and bidirectional ((V, T), (T, V)). As shown in LABEL:FF-SSM_input, single-order inputs exhibit significant performance variations: the (V, T) configuration achieves 83.1% mAP@0.5 (0.2% decrease from baseline), with particularly pronounced accuracy degradation for ’bicycle’ detection (1.0% drop). The (T, V) configuration shows more substantial performance decline, reaching only 82.5% mAP@0.5 (0.8% decrease) and suffering a 2.5% accuracy reduction for ’bicycle’ detection. These results reveal a critical feature attenuation phenomenon, where early input features are gradually forgotten during state propagation. To overcome this limitation, we propose a novel bidirectional architecture that establishes robust cross-modal feature memory pathways. Our solution not only elevates overall performance to 83.3% mAP@0.5 but also delivers significant improvements in ’bicycle’ detection accuracy (74.9%), while maintaining stable detection precision for both ’person’ and ’car’ categories.
As illustrated in Figure 9, we conducted a comprehensive comparison of three ERF map configurations for the FF-SSM module: (V, T), (T, V), and bidirectional ((V, T), (T, V)). The results demonstrate that our bidirectional input strategy yields a substantially expanded effective receptive field compared to unidirectional approaches. This empirical evidence strongly corroborates the efficacy of our proposed bidirectional architecture in enhancing the model’s receptive field capacity. The innovative design successfully mitigates feature attenuation in state-space models while maintaining linear computational complexity, thereby providing an efficient and effective solution for cross-modal feature fusion.
@@ 4.6 Qualitative Analysis
4.6.1 Statistical Analysis of the Model
Table 16: Statistical analysis of the model( Perform the experiment 10 times with random seeds released.).
Mean(%) VAR(%)
mAP@0.5 mAP mAP@0.5 mAP
FLIR 82.99 40.60 0.15 2.47
LLVIP 97.41 63.46 0.03 2.31
M3FD 89.11 59.76 0.18 0.42
The model demonstrates excellent robustness and stability based on the analysis of 10 random seed experiments as shown in LABEL:statiscal_analyse. On the FLIR dataset, it achieves an mAP@0.5 of 82.99% with a variance of only 0.15%. Its performance on the LLVIP dataset is particularly outstanding, reaching an mAP@0.5 of 97.41% with an extremely low variance of 0.03%. On the M3FD dataset, the model attains an mAP@0.5 of 89.11% and an mAP variance of just 0.42%. These results indicate that the model exhibits remarkable stability, fully demonstrating its strong resistance to variations in initial conditions. It maintains consistent performance across different data distributions, showcasing excellent practical engineering value.
@@ 4.6.2 Heatmap Visualization
Figure 10 presents the heatmap comparisons of three competing methods (baseline, ICAFusion and MS2Fusion) on RGB-T images. Through comprehensive analysis across multiple scenarios, the MS2Fusion method demonstrates consistent superiority. In the parking lot scenario (first row), the baseline method manages to detect partial instances of cars and pedestrians, yet produces significantly incomplete bounding boxes. Although ICAFusion shows noticeable improvement by capturing more objects, it still suffers from occasional missed detections. By contrast, MS2Fusion achieves near-perfect performance, precisely localizing all cars and pedestrians with highly accurate bounding boxes. The performance gap becomes even more pronounced in the challenging nighttime street scene (second row). While the baseline method can identify some pedestrians and vehicles, its detection accuracy proves inadequate. ICAFusion offers moderate improvement over the baseline, yet still struggles with low detection accuracy in these low-light conditions. Remarkably, MS2Fusion maintains excellent performance, reliably identifying all objects with exceptional precision even in this demanding scenario. Similarly, in the road scene (third row), the baseline method exhibits several limitations, including missed detections of vehicles and cyclists. ICAFusion partially addresses these issues by detecting more instances, but its accuracy remains suboptimal. Impressively, MS2Fusion again outperforms both competitors, achieving better detection performance with precise bounding boxes and minimal false negatives.